OpenStudy (anonymous):

t1 = 1; t2= 1; tn-1 = tn-1 - tn? (Recursion formula)

OpenStudy (amistre64):

its good to put subscripts in {curlys} to set it apart

OpenStudy (anonymous):

ohh.. okay

OpenStudy (amistre64):

t{1} = 1 t{2} = 1 t{n-1} = .....

OpenStudy (amistre64):

otherwise it just looks like normal mathing

OpenStudy (anonymous):

I tried: t{n - 1} = t{n-1} - t{n} t{2 - 1} = t3-1 - t2 t{1} = t{2} - t{2} t{1} = 1 - 1 t{1} = 0

OpenStudy (anonymous):

ohhh, i see.

OpenStudy (anonymous):

and the answers are: 1, 1, 0, -1, -1, 0

OpenStudy (amistre64):

t{n} = t{n-1} - t{n-1} is always 0, since you are subtracting the same amount from itself. it looks like a wierd recursion

OpenStudy (anonymous):

ohh, i see..

OpenStudy (amistre64):

is there a larger problem that this goes to?

OpenStudy (anonymous):

ohhhhh!!! Aww man, I wrote the question wrong all this time. It's: t{1} = 1; t{2} = 1; t{n+2} = t{n+1}-t{n}?

OpenStudy (amistre64):

that does look better ;)

OpenStudy (anonymous):

ohh, sorry D:

OpenStudy (amistre64):

t{1} = 1 t{2} = 1 t{n+2} = t{n+1}-t{n} t{3} = t{2} - t{1} = 0 t{4} = t{3} - t{2} = -1 t{5} = t{4} - t{3} = -1 t{6} = t{5} - t{4} = 0

OpenStudy (anonymous):

ooouu, thank you amistre64! :D